# The microscale organization of directed hypergraphs

### v1

@@ Results
To perform motif analysis on real-world directed hypergraphs, we propose an exact algorithm to count the frequency of all connected sub-hypergraph patterns and quantify their abundance with respect to a null model (see Methods). Given the intractability of the problem for large sub-hypergraphs, we limit our study of empirical data to patterns involving three and four nodes. Moreover, we focus on patterns that include at least one group interaction. In Figure 6, we show the most over-represented patterns of directed higher-order interactions with three and four nodes across different domains. Each domain reveals distinct motifs, characterized by different directed hyperedge types, sizes, densities and patterns of reciprocity. In terms of hyperedges types, E-mail and qna involve abundant patterns with only many-to-one and one-to-many interactions. Other datasets display more diverse patterns, including combinations of one-to-many, many-to-one, and many-to-many interactions (this is possible only in motifs with four nodes). Traditional one-to-one interactions are commonly part of abundant patterns in all datasets. The number of interactions in abundant sub-hypergraphs is small in the Metabolic and Citation domains, often involving just one or two hyperedges. In contrast, the E-mail, Bitcoin, and qna domains tend to be richer in interactions. This observation is reversed when considering the average size of interactions. This relation between the number and the average size of interactions aligns with previous studies on undirected higher-order motifs [19]. A common pattern in many datasets is the coexistence of group interactions, alongside lower-order interactions within the same set of nodes. These interactions seem to play a role in increasing the overall reciprocity of the patterns, suggesting the existence of a feedback mechanism. This is particularly evident in E-mail data. In addition to reciprocity, the direction of lower-order interactions in abundant patterns suggests a reinforcing mechanism, where subsets of source and target nodes interact at multiple interaction sizes. It is also likely to observe lower-order interactions among nodes that act together as a source or target of the group interaction.
@@ Discussion
In this work, we proposed new measures and tools to analyze the structural organization of directed hypergraphs at their microscale. First, we analyzed hyperedge signature vectors to identify the abundance of each hyperedge structure across datasets and identified classes of systems sharing similar higher-order connectivity patterns. Second, we introduced three distinct types of higher-order reciprocity measures: exact, strong, and weak reciprocity. Each definition offers a different perspective on how group interactions can be reciprocated, ranging from strict to more relaxed forms of reciprocal influence, and can be computed efficiently, making it suitable also for the analysis of very large systems. We showed that all systems exhibit reciprocity in broad terms, though different domains are associated with specific patterns and sensitivity to specific reciprocity measures. Lastly, we extended the notion of motifs to directed hypergraphs, capturing recurring patterns of directed interactions. Motif analysis revealed frequent microscale structures and highlighted common organizational principles playing a role in the function and behavior of systems, such as the existence of reinforcing or feedback mechanisms among dyadic and non-dyadic interactions in groups.
Taken together, by considering the nuances related to the directionality of interactions in directed hypergraphs, our research provides a framework to understand higher-order connectivity in directed complex systems, opening up a wide range of potential applications in diverse fields such as social network analysis, biology, and finance. For instance, the study of multi-party financial transactions as directed higher-order structures may capture more complex patterns of fraudulent activity than traditional graph-based models [45]. Similarly, directed hypergraphs may enhance the accuracy of existing frameworks in identifying and predicting important genes based on genomic expression relations [46]. As scalability is a pressing issue in hypergraph algorithms, future work may explore advanced techniques for detecting motifs in large-scale directed hypergraphs, including sampling methods [47], to expand our analysis beyond patterns of four nodes. Another interesting venue for further studies is related to the study of reciprocity in weighted or time-evolving hypergraphs, where interactions are associated with different intensities or specific moments in time. All in all, our work reveals new structural principles behind the organization of real-world systems, shedding light on the complex interplay between structural patterns and functionality in directed complex systems.
@@ Methods
In order to design efficient algorithms for mining directed higher-order motifs, we extend prior ideas developed for the same problem in undirected hypergraphs [47]. Our algorithms are efficient enough to count motifs of size $3$ and $4$ in datasets of reasonable size (comparable to those used in our experiments). However, scaling to larger datasets and motifs of larger size would require more sophisticated approaches, such as sampling algorithms [47], which we leave for future work. Further details on the execution times of the algorithms for mining motifs of orders $3$ and $4$ can be found in Supplementary Notes $2$.
The algorithm for mining motifs (involving at least one group interaction) of order $3$ begins by iterating through each hyperedge in the hypergraph that contains exactly three vertices. For each such hyperedge, it identifies all possible subsets of vertices and checks whether one or more subsets form valid directed hyperedges in the hypergraph. Valid subsets, along with the original hyperedge, define the motif structure involving those three vertices. To ensure consistency in motif identification, the algorithm generates a canonical form of the motif by lexicographical ordering its vertices and edges, which can be computed by sorting the $n!$ possible relabels. This canonical representation allows motifs with the same structural pattern to be compared and counted, even if they differ in their vertex labels. Each canonical form of motifs is stored in a frequency hash map. If the motif has not been encountered before, it is added to the map; if it has, its frequency count is incremented. In the end, the algorithm outputs a distribution of the various motif structures of order $3$. This algorithm operates in linear time with respect to the number of hyperedges of order $3$. Specifically, its computational complexity is $O(m_{3})$, where $m_{3}$ is the number of hyperedges involving exactly three vertices. Each motif construction and comparison is performed in constant time due to the fixed size of the motifs. For more details, refer to the pseudocode in Supplementary Note $2$.
The algorithm for mining motifs of order $4$ follows a similar approach. First, it iterates over all hyperedges of size $4$, counting the motifs involving exactly these $4$ nodes. Unlike the previous algorithm, it then iterates over all hyperedges of size $3$, performing an additional neighborhood exploration step to identify the fourth node involved in the motif. Each neighboring node is considered during this process. Once the 4 nodes are identified, the algorithm constructs the motif as before. The pseudocode for this algorithm is provided in Supplementary Note $2$.
@@
To distinguish meaningful, non-random interaction patterns from those that may occur by chance, we use a configuration model as a null model to evaluate the statistical significance of the interaction patterns after computing their frequency in our directed hypergraphs. The configuration model generates randomized versions of the original hypergraph while preserving key properties, such as the in-degree and out-degree sequences, as well as the source and target sizes of the hyperedges [35]. By comparing the observed frequencies with those found in the randomized networks, we can identify significantly over-represented motifs. In particular, each motif is associated with the abundance score $\Delta_{i}$ relative to random networks proposed in [44],
$\Delta_{i}=\frac{N\text{real}_{i}-\langle N\text{rand}_{i}\rangle}{N\text{real}_{i}+\langle N\text{rand}_{i}\rangle+\epsilon}$ (1)
Following [44, 19], we set $\epsilon=4$. We sample $N=10$ times from the configuration model.
@@
Available as part of Hypergraphx (HGX) [48].
@@
Data [39] is publicly available and also easily accessible through HGX [48].
@@ Acknowledgements
F.B. acknowledges support from the Air Force Office of Scientific Research under award number FA8655-22-1-7025. A.M. acknowledges support from the European Union through Horizon Europe CLOUDSTARS project (101086248).
@@
This section provides detailed descriptions of the datasets used in our experiments. The datasets are originally collected in [39] and represent a diverse range of real-world systems with directed higher-order interactions. Summary statistics of the datasets used in our experiments are reported in Table S1.
  • Question answering data. We use two qna datasets: Math-overflow and Server-fault, both sourced from Stack Exchange logs. A hyperedge $e_{i}=(S_{i},T_{i})$ indicates a question posted by the user in the target set $T_{i}$ and answered by the users in the source set $S_{i}$. Each hyperedge has a unit target set, i.e., $|T_{i}|=1,\forall i=\{1,...,|E|\}$.
  • Email data. We use two email datasets: email-enron [49] and email-eu [50]. A hyperedge $e_{i}=(S_{i},T_{i})$ represents an email where the sender is the source set $S_{i}$, and the receivers (including cc-ed users) form the target set $T_{i}$. Each hyperedge has a unit source set, i.e., $|S_{i}|=1,\forall i=\{1,...,|E|\}$.
  • Bitcoin transactions data. We use three bitcoin transaction datasets: bitcoin-2014, bitcoin-2015, and bitcoin-2016 [51]. They contain the first 1 500 00015000001\,500\,0001 500 000 transactions in 11/2014, 06/2015, and 01/2016 respectively. A hyperedge $e_{i}=(S_{i},T_{i})$ corresponds to a transaction where the accounts from which the coins are sent form the source set $S_{i}$, and the accounts receiving the coins make up the target set $T_{i}$.
  • Metabolic data. We use two methabolic datasets: iAF1260b and iJO1366 [52]. Nodes are the genes and hyperedges are metabolic reactions. A hyperedge $e_{i}=(S_{i},T_{i})$ indicates that the reaction among genes in the source set $S_{i}$ results in genes in the target set $T_{i}$.
  • Citation data. We use two citation datasets: citation-data mining and citation-software [53, 54]. A hyperedge $e_{i}=(S_{i},T_{i})$ represents a citation from a paper co-authored by the authors in the source set $S_{i}$ to a paper co-authored by the authors in the target set $T_{i}$. Papers with more than 10 authors are filtered out.
@@
Supplementary Note 2. Algorithms for motif analysis in directed hypergraphs
This section provides further details on the algorithms for motif analysis in directed hypergraphs. In Fig. S1, we show the execution times of our algorithms for motifs of order $3$ and $4$ across various datasets, highlighting the increase in time when moving from order $3$ to order $4$. More complex approaches will be needed to scale the analysis to larger motifs and larger dataset sizes. In Algorithm Supplementary Note 2. Algorithms for motif analysis in directed hypergraphs and Algorithm Supplementary Note 2. Algorithms for motif analysis in directed hypergraphs, we present detailed pseudocode for the algorithms designed to count directed higher-order motifs of sizes $3$ and $4$, respectively.
Figure S1: Execution times in seconds of the algorithms for mining motifs of order 3 and 4 across datasets. We consider $10$ trials for each dataset.
@@
[…]
@@
[…]
@@
17:end for each
$\mathcal{H}\leftarrow$Discard all hyperedges of order 4 from $\mathcal{H}$
@@
[…]

### v2

@@ Results
Given the intractability of the problem for large sub-hypergraphs, we limit our study of empirical data to patterns involving three and four nodes. Moreover, we focus on patterns that include at least one group interaction. Because the theoretical pattern space is still very large, while the empirical signal is sparse and concentrated (see Supplementary Note 3), we restrict our analysis to the positive tail of the motif z-score distributions (see Supplementary Note 3) to highlight structural differences across datasets. In particular, Figure 7 shows the most over-represented patterns of directed higher-order interactions with three and four nodes across different domains. Each domain reveals distinct motifs, characterized by different directed hyperedge types, sizes, densities and patterns of reciprocity. In terms of hyperedges types, E-mail and qna involve abundant patterns with only many-to-one and one-to-many interactions. Other datasets display more diverse patterns, including combinations of one-to-many, many-to-one, and many-to-many interactions (the latter is possible only in motifs with four nodes). Traditional one-to-one interactions are commonly part of abundant patterns in all datasets. The number of interactions in abundant sub-hypergraphs is small in the Bitcoin, Metabolic and Citation domains, often involving just one or two hyperedges. In contrast, the E-mail and qna domains tend to be richer in interactions. This observation is reversed when considering the average size of interactions. The relation between the number and the average size of interactions aligns with previous studies on undirected higher-order motifs [lotito2022higher]. A common pattern in many datasets is the coexistence of group interactions alongside lower-order interactions within the same set of nodes, a phenomenon often referred to as nestedness [lotito2022higher] or simpliciality [landry2024simpliciality]. Interestingly, considering the direction of such interactions reveals that they seem to play a role in increasing the overall reciprocity of the patterns, suggesting the existence of a feedback mechanism. This is particularly evident in E-mail data. In addition to reciprocity, the direction of lower-order interactions in abundant patterns suggests a reinforcing mechanism where subsets of source and target nodes interact at multiple interaction sizes. These observations are closely connected with the insights discussed in the previous sections about frequent co-senders and co-receivers nodes and higher-order reciprocity.
@@ Discussion
In this work, we proposed new measures and tools to analyze the structural organization of directed hypergraphs at their microscale. First, we analyzed hyperedge signature vectors to identify the abundance of each hyperedge structure across datasets and identified classes of systems sharing similar higher-order connectivity patterns. Second, we analyzed the excess overlap among source and target sets for each node in each system. The resulting distributions suggest that different domains may follow distinct organizational principles, ranging from redundant to more diverse interaction patterns. Then, we introduced three distinct types of higher-order reciprocity measures: exact, strong, and weak reciprocity. Each definition offers a different perspective on how group interactions can be reciprocated, ranging from strict to more relaxed forms of reciprocal influence, and can be computed efficiently, making it suitable also for the analysis of very large systems. We showed that all systems exhibit reciprocity in broad terms, though different domains are associated with specific patterns and sensitivity to specific reciprocity measures. Lastly, we extended the notion of motifs to directed hypergraphs, capturing recurring patterns of directed interactions. Motif analysis revealed frequent microscale structures and highlighted common organizational principles playing a role in the function and behavior of systems, such as the existence of reinforcing or feedback mechanisms among dyadic and non-dyadic interactions in groups.
Taken together, by considering the nuances related to the directionality of interactions in directed hypergraphs, our research provides a framework to understand higher-order connectivity in directed complex systems, opening up a wide range of potential applications in diverse fields such as social network analysis, biology, and finance. For instance, the study of multi-party financial transactions as directed higher-order structures may capture more complex patterns of fraudulent activity than traditional graph-based models [akoglu2015graph]. Similarly, directed hypergraphs may enhance the accuracy of existing frameworks in identifying and predicting important genes based on genomic expression relations [feng2021hypergraph]. As scalability is a pressing issue in hypergraph algorithms, future work may explore more advanced techniques for detecting motifs in large-scale directed hypergraphs. These include sampling methods already proposed for undirected higher-order motifs [lotito2024exact] or the use of parallel algorithms, which may achieve significant speed-ups beyond the implementation of our current Python library, thereby enabling analysis of motifs larger than four nodes. Another interesting venue for further studies is related to the study of reciprocity in time-evolving hypergraphs, since it can affect mechanisms of group formation [cencetti2021temporal, iacopini2024temporal] and inform the efficient seeding of information [mancastroppa2023hyper, genetti2024influence]. All in all, our work reveals new structural principles behind the organization of real-world systems, shedding light on the complex interplay between structural patterns and functionality in directed complex systems.
@@ Methods
Microcanonical set-swap configuration model for directed hypergraphs
We generate randomized counterparts of directed hypergraphs using a microcanonical set-swap configuration model. This approach is similar to the one proposed by Preti et al. [preti2024higher], and more recently by Kraakman et al. [kraakman2025hypercurveball]. More broadly, our null model fits within the family of configuration- and entropy-based random hypergraph ensembles that preserve degree and edge-size sequences [chodrow2020configuration, barthelemy2022class, nakajima2021randomizing, saracco2025entropy], which have been used in applications such as extracting statistically validated higher-order interactions [musciotto2021detecting] and motifs [lotito2022higher]. In our framework, each directed hyperedge $e\in E$ is specified by two disjoint node sets: a source set $s(e)\subseteq V$ and a target set $t(e)\subseteq V$, with $s(e)\cap t(e)=\varnothing$. The microcanonical ensemble comprises all simple directed hypergraphs that match the observed node out-/in-degree sequences (counts of appearances in $s(\cdot)$ and $t(\cdot)$, respectively) and preserve every hyperedge’s cardinalities $|s(e)|$ and $|t(e)|$. We forbid duplicated hyperedges.
Starting from the observed $H$, we repeatedly attempt set-swaps on one side at a time. In a source-side attempt, we select two distinct hyperedges $e_{i}\neq e_{j}$ uniformly and sample $u\in s(e_{i})$ and $v\in s(e_{j})$. We propose
$s^{\prime}(e_{i})=(s(e_{i})\setminus\{u\})\cup\{v\},\qquad s^{\prime}(e_{j})=(s(e_{j})\setminus\{v\})\cup\{u\},$
leaving $t(e_{i})$ and $t(e_{j})$ unchanged. The move is accepted only if it respects set semantics on the chosen side (no duplicates), preserves within-edge disjointness ($s(e)\cap t(e)=\varnothing$), and does not create a duplicate $(s^{\prime}(e),t(e))$. Target-side attempts are defined symmetrically. Accepted swaps conserve each node’s in/out degree and all $|s(e)|,|t(e)|$. Rejected proposals leave the state unchanged.
Hyperedge overlap
For each node, we quantify its hyperedge overlap separately for its participation in target sets (in-hyperedges) and source sets (out-hyperedges). For in-hyperedge overlap, we consider all hyperedges $e$ for which the node is in the target set $t(e)$. Let $\mathcal{E}_{\text{in}}$ be the collection of these hyperedges. We define the in-hyperedge overlap as
$O_{\text{in}}=\frac{\sum_{e\in\mathcal{E}_{\text{in}}}|t(e)|}{|\mathcal{E}_{\text{in}}|\left|\bigcup_{e\in\mathcal{E}_{\text{in}}}t(e)\right|},$
where $|t(e)|$ denotes the number of nodes in the target set of hyperedge $e$.
Similarly, for out-hyperedge overlap, we consider only hyperedges $e$ in which the node is a source (i.e., belongs to $s(e)$). Let $\mathcal{E}_{\text{out}}$ denote this collection, and define
$O_{\text{out}}=\frac{\sum_{e\in\mathcal{E}_{\text{out}}}|s(e)|}{|\mathcal{E}_{\text{out}}|\left|\bigcup_{e\in\mathcal{E}_{\text{out}}}s(e)\right|},$
with $|s(e)|$ being the size of the source set of hyperedge $e$. These metrics yield a value of 1 when all corresponding hyperedges share an identical set of nodes (i.e., maximal overlap), and decrease as the sets become more diverse.
To assess statistical significance, we compare the observed overlaps to those computed on an ensemble of randomized networks that preserve key structural properties (e.g., node degrees and hyperedge sizes). For each node $i$, we compute a z-score that standardizes the observed overlap relative to the null ensemble:
$z_{O}(i)=\frac{O_{i}^{\mathrm{obs}}-\langle O_{i}^{\mathrm{rand}}\rangle}{\sqrt{\operatorname{Var}\!\left(O_{i}^{\mathrm{rand}}\right)}}\,.$
We compute this for both $O_{\text{sources}}$ and $O_{\text{targets}}$.
@@
In order to design efficient algorithms for mining directed higher-order motifs, we extend prior ideas developed for the same problem in undirected hypergraphs [lotito2024exact]. Our algorithms are efficient enough to count motifs of size $3$ and $4$ in datasets of reasonable size (comparable to those used in our experiments). However, scaling to larger datasets and motifs of larger size would require more sophisticated approaches, such as sampling algorithms [lotito2024exact], which we leave for future work. Further details on the execution times of the algorithms for mining motifs of orders $3$ and $4$ can be found in Supplementary Notes $2$.
The algorithm for mining motifs (involving at least one group interaction) of order $3$ begins by iterating through each hyperedge in the hypergraph that contains exactly three vertices. For each such hyperedge, it identifies all possible subsets of vertices and checks whether one or more subsets form valid directed hyperedges in the hypergraph. Valid subsets, along with the original hyperedge, define the motif structure involving those three vertices. To ensure consistency in motif identification, the algorithm generates a canonical form of the motif by lexicographical ordering its vertices and edges, which can be computed by sorting the $n!$ possible relabels. This canonical representation allows motifs with the same structural pattern to be compared and counted, even if they differ in their vertex labels. Each canonical form of motifs is stored in a frequency hash map. If the motif has not been encountered before, it is added to the map; if it has, its frequency count is incremented. In the end, the algorithm outputs a distribution of the various motif structures of order $3$. This algorithm operates in linear time with respect to the number of hyperedges of order $3$. Specifically, its computational complexity is $O(m_{3})$, where $m_{3}$ is the number of hyperedges involving exactly three vertices. Each motif construction and comparison is performed in constant time due to the fixed size of the motifs. For more details, refer to the pseudocode in Supplementary Note $2$. The algorithm for mining motifs of order $4$ follows a similar approach. First, it iterates over all hyperedges of size $4$, counting the motifs involving exactly these $4$ nodes. Unlike the previous algorithm, it then iterates over all hyperedges of size $3$, performing an additional neighborhood exploration step to identify the fourth node involved in the motif. Each neighboring node is considered during this process. Once the 4 nodes are identified, the algorithm constructs the motif as before. The pseudocode for this algorithm is provided in Supplementary Note $2$.
@@
To distinguish meaningful, non-random interaction patterns from those that may occur by chance, we use a configuration model as a null model to evaluate the statistical significance of the interaction patterns after computing their frequency in our directed hypergraphs. The configuration model generates randomized versions of the original hypergraph while preserving key properties, such as the in-degree and out-degree sequences, as well as the source and target sizes of the hyperedges [preti2024higher]. By comparing the observed frequencies with those found in the randomized networks, we can identify significantly over-represented motifs.
In particular, each motif $i$ is associated with a standardized score $z_{M}(i)$, which quantifies how many standard deviations the observed frequency of the motif differs from its expected value under the configuration model [milo2002network, lotito2022higher]. A larger absolute value of $z_{M}(i)$ indicates a stronger statistical deviation from the null expectation, meaning the motif is significantly over- or under-represented compared to random networks. The score is defined as:
$z_{M}(i)=\frac{N_{i}^{\mathrm{obs}}-\langle N_{i}^{\mathrm{rand}}\rangle}{\sqrt{\operatorname{Var}\!\left(N_{i}^{\mathrm{rand}}\right)}}\,.$
Here, $N_{i}^{\mathrm{obs}}$ denotes the observed frequency of motif $i$ in the empirical hypergraph, and $\langle N_{i}^{\mathrm{rand}}\rangle$ and $\operatorname{Var}\!\left(N_{i}^{\mathrm{rand}}\right)$ are the mean and variance of motif $i$’s frequency across random hypergraphs generated by the configuration model. Following [milo2004superfamilies], we generate 100 random samples from the configuration model for each hypergraph to estimate the mean and variance of the motif frequencies.
@@
The tools for the analysis of directed hypergraphs presented in this work are available as part of Hypergraphx (HGX) [lotito2023hypergraphx].
@@
Data [kim2023reciprocity] is publicly available and also easily accessible through HGX [lotito2023hypergraphx].
@@ Acknowledgements
F.B. acknowledges support from the Air Force Office of Scientific Research under award number FA8655-22-1-7025. F.B. acknowledges support from the Austrian Science Fund (FWF) through project 10.55776/PAT1052824 and project 10.55776/PAT1652425. A.M. acknowledges support from the European Union through Horizon Europe CLOUDSTARS project (101086248).
@@
This section provides detailed descriptions of the datasets used in our experiments. The datasets are originally collected in [kim2023reciprocity] and represent a diverse range of real-world systems with directed higher-order interactions. Summary statistics of the datasets used in our experiments are reported in Table S1.
  • Question answering data. We use two qna datasets: Math-overflow and Server-fault, both sourced from Stack Exchange logs. A hyperedge $e_{i}=(S_{i},T_{i})$ indicates a question posted by the user in the target set $T_{i}$ and answered by the users in the source set $S_{i}$. Each hyperedge has a unit target set, i.e., $|T_{i}|=1,\forall i=\{1,...,|E|\}$. If the node in the target set also appears in the source set, we remove it from the source set.
  • Email data. We use two email datasets: email-enron [chodrow2020annotated] and email-eu [jure2014snap]. A hyperedge $e_{i}=(S_{i},T_{i})$ represents an email where the sender is the source set $S_{i}$, and the receivers (including cc-ed users) form the target set $T_{i}$. Each hyperedge has a unit source set, i.e., $|S_{i}|=1,\forall i=\{1,...,|E|\}$. If the node in the source set also appears in the target set, we remove it from the target set.
  • Bitcoin transactions data. We use three bitcoin transaction datasets: bitcoin-2014, bitcoin-2015, and bitcoin-2016 [wu2021detecting]. They contain the first 1 500 0001\,500\,000 transactions in 11/2014, 06/2015, and 01/2016 respectively. A hyperedge $e_{i}=(S_{i},T_{i})$ corresponds to a transaction where the accounts from which the coins are sent form the source set $S_{i}$, and the accounts receiving the coins make up the target set $T_{i}$.
  • Metabolic data. We use two methabolic datasets: iAF1260b and iJO1366 [yadati2020nhp]. Nodes are the genes and hyperedges are metabolic reactions. A hyperedge $e_{i}=(S_{i},T_{i})$ indicates that the reaction among genes in the source set $S_{i}$ results in genes in the target set $T_{i}$.
  • Citation data. We use two citation datasets: citation-data mining and citation-software [sinha2015overview, yadati2021graph]. A hyperedge $e_{i}=(S_{i},T_{i})$ represents a citation from a paper co-authored by the authors in the source set $S_{i}$ to a paper co-authored by the authors in the target set $T_{i}$. Papers with more than 10 authors are filtered out.
@@
Supplementary Note 2. Null model
Algorithm 1 Microcanonical set-swap configuration model
1:Directed hypergraph $H=(V,E)$ with source sets $s(e)$ and target sets $t(e)$; iteration budget $T$
[…]
5:for $t\leftarrow 1$ to $T$ do
6:  choose side from {source, target}
7:  sample distinct edges $e_{i}\neq e_{j}$ uniformly from $E$
8:  if side = source then
9:   pick $u\in s(e_{i})$, $v\in s(e_{j})$
10:   $s^{\prime}(e_{i})\leftarrow(s(e_{i})\setminus\{u\})\cup\{v\}$
11:   $s^{\prime}(e_{j})\leftarrow(s(e_{j})\setminus\{v\})\cup\{u\}$
12:   if $v\notin s(e_{i})$ and $u\notin s(e_{j})$ and $v\notin t(e_{i})$ and $u\notin t(e_{j})$ and $(s^{\prime}(e_{i}),t(e_{i}))\notin\Sigma$ and $(s^{\prime}(e_{j}),t(e_{j}))\notin\Sigma$ then
13:     accept: $s(e_{i})\leftarrow s^{\prime}(e_{i})$; $s(e_{j})\leftarrow s^{\prime}(e_{j})$; update $\Sigma$
14:   end if
15:  else
16:   pick $u\in t(e_{i})$, $v\in t(e_{j})$
17:   $t^{\prime}(e_{i})\leftarrow(t(e_{i})\setminus\{u\})\cup\{v\}$
18:   $t^{\prime}(e_{j})\leftarrow(t(e_{j})\setminus\{v\})\cup\{u\}$
19:   if $v\notin t(e_{i})$ and $u\notin t(e_{j})$ and $v\notin s(e_{i})$ and $u\notin s(e_{j})$ and $(s(e_{i}),t^{\prime}(e_{i}))\notin\Sigma$ and $(s(e_{j}),t^{\prime}(e_{j}))\notin\Sigma$ then
20:     accept: $t(e_{i})\leftarrow t^{\prime}(e_{i})$; $t(e_{j})\leftarrow t^{\prime}(e_{j})$; update $\Sigma$
21:   end if
22:  end if
23:end for
24:return $H^{\prime}$;
Supplementary Note 3. Additional results about motif analysis in directed hypergraphs
The combinatorial explosion in the number of possible subhypergraph patterns for a given number of nodes makes it impractical to report statistics for every single motif. In the main text, we therefore only display the visual representations of the most over-represented patterns. Here, we expand this analysis and provide additional quantitative results on the abundance and significance of motifs across datasets.
Figure S1 provides an overview of the structural diversity and frequency distribution of motifs. Panel (a) reports the number of distinct non-isomorphic subhypergraph patterns with three and four nodes observed across datasets. Although the total number of possible non-isomorphic configurations grows superexponentially with the number of nodes, only a subset of them actually appear in real data. Panel (b) examines how motif occurrences are distributed among these observed patterns, showing whether they are concentrated in a few motifs or spread more evenly. The results indicate that the top 10% of motifs account for most occurrences. Some variability is visible across domains; for instance, 3-node motifs in metabolic and email datasets display a slightly more balanced distribution. Overall, motif frequencies are highly concentrated in a relatively small subset of distinct patterns.
Figure S1: (a) Number of observed non-isomorphic subhypergraph patterns with three and four nodes in each dataset. (b) Cumulative fraction of motif occurrences versus the ranked motif index, showing that few motifs account for most occurrences.
To assess the statistical significance of these motifs, we compute the z-scores (see Methods) and analyze their distribution, shown in Fig. S2. The distributions, capped between 10-10 and $10$, reveal the range of significance levels across datasets. The motif drawings presented in the main paper (see Fig. $7$) correspond to those lying in the positive tail of the distributions, that is, motifs whose observed frequencies deviate by several standard deviations from the null expectation.
Figure S2: Distribution of motif z-scores across datasets, capped between 10-10 and $10$. Motifs shown in the main paper correspond to those in the positive tail of the distribution, indicating strong over-representation relative to the configuration model.
Supplementary Note 4. Algorithms for motif analysis in directed hypergraphs
This section provides further details on the algorithms for motif analysis in directed hypergraphs. In Table S2, we report the execution times of our algorithms for motifs of order $3$ and $4$ across various datasets, highlighting the increase in time when moving from order $3$ to order $4$. More complex approaches will be needed to scale the analysis to larger motifs and larger dataset sizes. In Algorithm Supplementary Note 4. Algorithms for motif analysis in directed hypergraphs and Algorithm Supplementary Note 4. Algorithms for motif analysis in directed hypergraphs, we present detailed pseudocode for the algorithms designed to count directed higher-order motifs of sizes $3$ and $4$, respectively.
Table S2: Execution times by dataset for Order 3 and Order 4 (mean and standard deviation, in seconds).
Order 3 (s) Order 4 (s)
Dataset Mean $\pm$ Std Mean $\pm$ Std
[…]
email-enron $3±\pm 0.000$$3.27±\pm 0.0$30.03
[…]
metabolic-iaf1260b $1±\pm 0.000$.00 $24±\pm 0.0$00.00
metabolic-ijo1366 $1±\pm 0.000$.00 $27±\pm 0.0$10.01
[…]
@@
11:  $C_{m}\leftarrow$ lexicographically minimum canonic relabel of $\mathit{motif}$
@@
11:  $C_{m}\leftarrow$ lexicographically minimum canonic relabel of $\mathit{motif}$
@@
30:     $C_{m}\leftarrow$ lexicographically minimum canonic relabel of $\mathit{motif}$
@@
Supplementary Note 5. Graphical outlines of the algorithms proposed
Here, we provide a schematic flowchart representation of all algorithmic components introduced in the main text. The diagrams summarize the logical structure, decision steps, and computational complexity of each procedure. They are intended to complement the formal descriptions and pseudocode, highlighting the conceptual flow of the algorithms used for the set-swap configuration model (Fig. S3), reciprocity evaluation (Fig. S4) and motif discovery (Fig. S5).
Figure S3: Microcanonical set-swap configuration model. Starting from an observed hypergraph $H=(V,E)$, two edges are randomly selected at each iteration, and one vertex is swapped between their source or target sets. The swap is accepted only if it does not produce duplicates. Repeating this process $T$ times preserves node in/out-degrees and edge sizes, generating randomized but structurally consistent hypergraphs.
Figure S4: Reciprocity algorithms. On the left, the exact reciprocity algorithm stores each hyperedge and checks for its reverse. In the middle, the strong reciprocity algorithm accumulates nodes reachable from the target set to verify full coverage of the source set. On the right, the weak reciprocity algorithm builds all source–target pairs and checks for any reverse target–source connection.
Figure S5: Motif discovery algorithms. On the left, order-3 motifs are identified from each 3-vertex hyperedge by collecting and relabelling connected subhyperedges. In the middle, order-4 motifs are extracted from single 4-vertex hyperedges using the same procedure. On the right, additional order-4 motifs arise from pairs of adjacent 3-vertex hyperedges spanning four vertices.